### v1

# arXiv:2602.23440v1[cs.CL] 26 Feb 2026

Abstract
Training large language models to reason with search engines via reinforcement learning is hindered by a fundamental credit assignment problem: existing methods such as Search-R1 provide only a sparse outcome reward after an entire multi-step trajectory, making it infeasible to attribute success or failure to individual reasoning and retrieval decisions. Process-reward methods like StepSearch alleviate this by introducing step-level supervision, but rely on heuristic rewards such as TF-IDF overlap with gold documents, and still sample $k$ complete trajectories per example, retaining high gradient variance. We propose Slate, a framework built on two complementary ideas: (1) truncated step-level sampling, which generates $k$ trajectories that share a common prefix and differ only at the next step, and (2) dense LLM-as-judge rewards, which replace heuristic scoring with a capable LLM evaluator that assesses the quality of each reasoning step, search query, and answer, providing richer and more reliable supervision. We theoretically prove that under the same dense reward structure, truncated sampling reduces the variance of advantage estimates by up to a factor of $T$ compared to full-trajectory sampling for $T$-step trajectories, yielding lower-variance, better-targeted policy gradients. Experiments on seven QA benchmarks confirm that Slate consistently outperforms both sparse-reward and process-reward baselines, with the largest gains on harder multi-hop tasks and smaller models.
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation (Hendrycks et al., 2020; Clark et al., 2018). Despite these achievements, LLMs often struggle with complex reasoning tasks (Wei et al., 2022) and lack access to up-to-date external knowledge (Jin et al., 2024). Integrating search engines into the LLM reasoning loop, where the model interleaves its own chain-of-thought reasoning with external retrieval calls, has emerged as a promising paradigm for knowledge-intensive question answering (Yao et al., 2023; Trivedi et al., 2022a).
Reinforcement learning (RL) provides a natural framework for optimizing such search-augmented reasoning systems. Search-R1 (Jin et al., 2025) pioneered the use of RL (PPO and GRPO) to train LLMs that autonomously invoke search engines during multi-turn reasoning, using outcome-based exact-match rewards. While effective, this sparse reward design suffers from a fundamental limitation: the credit assignment problem. When the model receives a single binary signal after completing a multi-step trajectory, potentially involving several rounds of thinking, query generation, and information processing, it cannot determine which intermediate steps contributed positively or negatively to the outcome. Recent work has attempted to address this limitation through process-level supervision. StepSearch (Wang et al., 2025) augments PPO with step-wise rewards based on information gain and redundancy penalties, achieving notable improvements on multi-hop QA benchmarks. However, StepSearch still relies on sampling $k$ complete trajectories per training example, which is computationally expensive and retains variance from all steps in the gradient estimates for any individual step.
In this paper, we propose Slate111code available at https://github.com/algoprog/SLATE, Step-Level Advantage estimation for Truncated Exploration, which addresses these limitations through two complementary ideas:
  • Truncated Step-Level Sampling: Instead of sampling $k$ full independent trajectories, we sample $k$ truncated trajectories that share a common prefix $\tau_{<t}$ and differ only at step $t$. This allows GRPO-style group relative advantages to be computed at the step level, directly attributing rewards to the specific action that caused them.
  • Dense LLM-as-Judge Rewards: We replace the sparse EM outcome reward with dense, step-level rewards produced by an LLM evaluator. The judge scores each reasoning step (thinking quality), each search query (query quality), and the final answer (correctness) on a discrete scale $\{-1,0,+1\}$, providing rich supervision at every decision point.
We provide formal theoretical analysis showing that truncated step-level sampling yields advantage estimates with provably lower variance than standard full-trajectory sampling, with up to a $T$-fold reduction for $T$-step trajectories, resulting in lower-variance policy gradients. Combined with the dense reward signal from the LLM judge, this enables faster convergence, better credit assignment, and improved final performance. Our contributions are the following:
  1. 1.
    We propose truncated step-level sampling for GRPO, an exploration strategy that generates $k$ alternative continuations only at the current decision point rather than $k$ complete trajectories. We formally prove this achieves a $T$-fold advantage variance reduction over full-trajectory sampling, yielding lower-variance policy gradients.
  2. 2.
    We introduce an LLM-as-judge reward system that provides dense, discrete step-level supervision for reasoning quality, search query quality, and answer correctness, replacing the sparse outcome reward.
  3. 3.
    We demonstrate that Slate significantly outperforms both existing sparse-reward methods (Search-R1) and process-reward methods (StepSearch) on seven QA benchmarks.
2 Related Work
LLMs and Retrieval
Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) integrates external knowledge into LLM generation by retrieving relevant passages based on the input query. While effective for simple lookups, standard RAG struggles with multi-hop reasoning where iterative retrieval is needed (Yang et al., 2018; Trivedi et al., 2022a). Alternative approaches treat the search engine as a tool (Yao et al., 2023; Schick et al., 2023), enabling the LLM to decide when and what to search. However, prompting-based tool-use methods rely on in-context examples and fail to generalize robustly, while supervised fine-tuning approaches require expensive annotated trajectories (Schick et al., 2023; Asai et al., 2024).
RL for LLM Reasoning
Reinforcement learning has emerged as a potent paradigm for enhancing LLM reasoning, as demonstrated by OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). Policy gradient methods including PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) have been widely adopted for training reasoning models. GRPO is particularly attractive as it eliminates the need for a separate critic model by using group-relative advantages computed from multiple sampled responses. However, applying these methods to search-augmented scenarios introduces unique challenges around reward design, credit assignment, and integration of retrieved content.
Search-Augmented RL
Search-R1 (Jin et al., 2025) introduced the paradigm of RL training with interleaved search engine calls, using outcome-based EM rewards and retrieved token loss masking. Follow-up work includes R1-Searcher (Song et al., 2025), ReSearch (Chen et al., 2025), and ZeroSearch (Sun et al., 2025), all of which rely on sparse global rewards. StepSearch (Wang et al., 2025) addresses the reward sparsity problem by introducing step-wise rewards with information gain and redundancy penalties, but still samples complete trajectories and relies on access to ground-truth intermediate documents for computing information gain. Our work differs from all prior methods by combining truncated step-level sampling with model-based dense rewards from an LLM judge (eliminating the need for ground-truth intermediate annotations).
Figure 1: Comparison of GRPO (with full trajectory sampling) and our truncated step-level sampling. By fixing the prefix $\tau_{<t}$, all variation in the sampled group is localized to step $t$.
3 Methodology
We present Slate, a training framework for search-augmented LLM reasoning that combines truncated step-level sampling with dense LLM-as-judge rewards. We build on the multi-turn search interaction framework of Search-R1 (Jin et al., 2025) and optimize using a modified GRPO objective. Figure 1 provides a high-level overview.
3.1 Preliminaries: Search-Augmented RL
Following Search-R1, we model the search engine $\mathcal{E}$ as part of the environment. The LLM policy $\pi_{\theta}$ generates outputs interleaved with search engine calls, producing trajectories of the form:
$\tau=\underbrace{\langle\texttt{think}\rangle s_{1}\langle\texttt{/think}\rangle}_{\text{reasoning}}\;\underbrace{\langle\texttt{search}\rangle q_{1}\langle\texttt{/search}\rangle}_{\text{query}}\;\underbrace{\langle\texttt{info}\rangle d_{1}\langle\texttt{/info}\rangle}_{\text{retrieval}}\;\cdots$ (1)
concluding with $\langle\texttt{answer}\rangle a\langle\texttt{/answer}\rangle$, where $s_{t}$ denotes the reasoning at step $t$, $q_{t}$ the search query, $d_{t}=\mathcal{E}(q_{t})$ the retrieved documents, and $a$ the final answer. We denote the number of search steps as $T$. The standard RL objective with search is:
$\max_{\pi_{\theta}}\;$ $\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi_{\theta}(\cdot\mid x;\mathcal{E})}\left[r_{\phi}(x,y)\right]-\beta\,\mathbb{D}_{\text{KL}}\!\left[\pi_{\theta}(y\mid x;\mathcal{E})\,\|\,\pi_{\text{ref}}(y\mid x;\mathcal{E})\right],$
where $r_{\phi}(x,y)$ is the reward function and $\pi_{\text{ref}}$ is the reference policy.
Standard GRPO.
In GRPO (Shao et al., 2024), $G$ complete trajectories $\{y_{1},\ldots,y_{G}\}$ are sampled for each input $x$, and the advantage for trajectory $i$ is:
$\hat{A}_{i}=\frac{R(y_{i})-\mu_{R}}{\sigma_{R}+\epsilon},$ (2)
where $\mu_{R}=\frac{1}{G}\sum_{i=1}^{G}R(y_{i})$, $\sigma_{R}=\sqrt{\frac{1}{G}\sum_{i=1}^{G}(R(y_{i})-\mu_{R})^{2}}$, and $R(y_{i})$ is the trajectory-level reward. As discussed in Section 1, this suffers from poor credit assignment (a single scalar weights gradients for all $T$ steps) and high variance ($\hat{A}_{i}$ reflects variation across all steps).
3.2 Truncated Step-Level Sampling
Our key algorithmic novelty is truncated step-level sampling: instead of sampling $k$ complete independent trajectories that may diverge from the very first step, we sample $k$ truncated trajectories that share a common prefix and differ only at the next reasoning step $t$.
Formal Definition.
Let $\tau_{<t}=(s_{1},q_{1},d_{1},\ldots,s_{t-1},q_{t-1},d_{t-1})$ denote the trajectory prefix up to (but not including) step $t$. At each step $t$, we generate $k$ candidate next-step actions by sampling from the policy conditioned on the shared prefix:
$a_{t}^{(j)}=(s_{t}^{(j)},q_{t}^{(j)})\sim\pi_{\theta}\!\left(\cdot\mid x,\tau_{<t}\right),\quad j=1,\ldots,k.$ (3)
Here, each $a_{t}^{(j)}$ consists of a reasoning step $s_{t}^{(j)}$ (the <think> block) followed by a search query $q_{t}^{(j)}$ (the <search> block), or alternatively a final answer $a^{(j)}$ (the <answer> block) if the model chooses to terminate. Each candidate action $a_{t}^{(j)}$ is then evaluated by the LLM-as-judge reward model (Section 3.3) to obtain a step-level reward $r_{t}^{(j)}$. The step-level group-relative advantage is:
$\hat{A}_{t}^{(j)}=\frac{r_{t}^{(j)}-\bar{r}_{t}}{\sigma_{t}+\epsilon},$ (4)
where $\bar{r}_{t}$ and $\sigma_{t}$ are the mean and standard deviation of rewards within the step-$t$ group.
Trajectory Construction.
After computing advantages for all $k$ candidates at step $t$, we select the action to continue the trajectory. The selected action $a_{t}^{(j^{*})}$ is appended to the prefix, the search engine is invoked to retrieve documents $d_{t}=\mathcal{E}(q_{t}^{(j^{*})})$, and the process repeats at step $t+1$. The selection can follow different strategies:
  • Best-of-$k$: $j^{*}=\arg\max_{j}r_{t}^{(j)}$ (pure exploitation).
  • Reward-weighted sampling: $j^{*}\sim\text{softmax}(\hat{A}_{t}^{(1)},\ldots,\hat{A}_{t}^{(k)}\,/\,\eta)$ with temperature $\eta$ (exploration-exploitation trade-off).
In our experiments we adopt reward-weighted sampling (with temperature $\eta$), as it balances exploitation of high-reward actions with exploration of diverse reasoning paths, preventing the trajectory from collapsing to a single greedy mode early in training. The complete procedure is presented in Algorithm 1 (Appendix).
3.3 Dense LLM-as-Judge Rewards
We replace the sparse outcome reward of Search-R1 with dense, step-level rewards produced by an LLM evaluator. At each step $t$, the reward model evaluates the quality of (1) the reasoning step, (2) the search query, and (3) the final answer (at the last step). All rewards use a discrete ternary scale $\{-1,0,+1\}$. Crucially, the LLM judge is prompted to produce a chain-of-thought reasoning before outputting the final score; we found that this “reason-then-score” protocol substantially improves the reliability and consistency of the reward signal compared to directly predicting a score.
Thinking Reward.
The thinking reward $r_{\text{think}}(s_{t},\tau_{<t})\in\{-1,0,+1\}$ evaluates the quality of reasoning step $s_{t}$ given the trajectory context $\tau_{<t}$. The LLM judge assesses five criteria: relevance, clarity, specificity, progress, and faithfulness. A score of +1+1 indicates clear, relevant reasoning that identifies specific information needs; 0 indicates somewhat relevant but vague reasoning; and 1-1 indicates irrelevant, misleading, or counterproductive reasoning.
Query Reward.
The query reward $r_{\text{query}}(q_{t},s_{t},\tau_{<t})\in\{-1,0,+1\}$ evaluates the search query $q_{t}$ conditioned on the retrieved documents, the preceding reasoning and the trajectory context. The judge assesses five criteria: relevance, specificity, searchability, alignment, and novelty. A score of +1+1 indicates a specific, well-formed query with clear keywords; 0 indicates some specificity but generic phrasing; and 1-1 indicates an irrelevant or poorly formed query.
Final Answer Reward.
The answer reward $r_{\text{answer}}(a,a_{\text{gold}},\tau)\in\{-1,0,+1\}$ evaluates whether the predicted answer $a$ conveys the same information as the ground truth $a_{\text{gold}}$: +1+1 for correct, 0 for partially correct or ambiguous, and 1-1 for incorrect. Unlike the binary EM reward in Search-R1, this ternary signal distinguishes between partially correct and fully incorrect answers, and can handle paraphrases and semantic equivalence.
3.3.1 Composite Step Reward
The total reward for action $a_{t}^{(j)}=(s_{t}^{(j)},q_{t}^{(j)})$ at step $t$ is:
$r_{t}^{(j)}=r_{\text{think}}(s_{t}^{(j)},\tau_{<t})+r_{\text{query}}(q_{t}^{(j)},s_{t}^{(j)},\tau_{<t}).$ (5)
When the model produces an answer at step $t$, the reward additionally includes the answer component and an early-termination bonus:
$r_{t}^{(j)}=r_{\text{think}}(s_{t}^{(j)},\tau_{<t})+r_{\text{answer}}(a^{(j)},a_{\text{gold}},\tau)+\lambda\cdot\frac{B-t}{B},$ (6)
where $B$ is the maximum action budget and $\lambda\geq 0$ controls the strength of the early-termination bonus.
Early-Termination Bonus.
The bonus term $\lambda\cdot(B-t)/B$ in Eq. 6 encourages the model to answer as soon as it has gathered sufficient information, rather than issuing superfluous search queries. Since the bonus only applies to answer candidates, it creates a meaningful advantage signal within each step-level group (Appendix A.8). Unlike StepSearch (Wang et al., 2025), which requires ground-truth documents, our reward requires only the trajectory context and gold final answer.
3.4 Step-Level GRPO Optimization
We now describe how the truncated step-level samples and dense rewards are integrated into the GRPO optimization framework.
Step-Level Policy Gradient.
At each step $t$, given $k$ candidate actions $\{a_{t}^{(1)},\ldots,a_{t}^{(k)}\}$ with step-level advantages $\{\hat{A}_{t}^{(1)},\ldots,\hat{A}_{t}^{(k)}\}$ (Eq. 4), we compute the clipped policy gradient objective for each candidate. Following Search-R1, we apply loss masking to retrieved tokens: let $I(y_{l})=1$ if $y_{l}$ is generated by the LLM and $I(y_{l})=0$ if $y_{l}$ is a retrieved token. The step-level objective is:
$\mathcal{J}_{t}^{(j)}(\theta)=\frac{1}{\sum_{l}I(y_{l})}\sum_{\begin{subarray}{c}l:\,I(y_{l})=1\end{subarray}}\min\big(\rho_{l}\,\hat{A}_{t}^{(j)},\text{clip}(\rho_{l},1{-}\epsilon,1{+}\epsilon)\,\hat{A}_{t}^{(j)}\big),$ (7)
where $\rho_{l}=\pi_{\theta}(y_{l}\mid x,y_{<l};\mathcal{E})/\pi_{\theta_{\text{old}}}(y_{l}\mid x,y_{<l};\mathcal{E})$ is the per-token importance ratio and the summation runs only over LLM-generated tokens in action $a_{t}^{(j)}$. The complete Slate training objective aggregates over all steps and all candidates:
$\mathcal{J}_{\text{{Slate}{}}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\bigg[\sum_{t=1}^{T}\frac{1}{k}\sum_{j=1}^{k}\mathcal{J}_{t}^{(j)}(\theta)-\beta\,\mathbb{D}_{\text{KL}}\!\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\bigg],$
where $\beta$ is the KL regularization coefficient and the KL divergence is computed only over LLM-generated tokens.
4 Theoretical Analysis
We formally analyze the variance reduction from truncated step-level sampling. Consider a $T$-step trajectory $\tau=(a_{1},\ldots,a_{T})$ with step-level rewards $r_{t}=r(a_{t},\tau_{<t})$. To isolate the effect of the sampling strategy, we compare two estimators under the same additive reward $R(\tau)=\sum_{t}r_{t}$: (A) standard GRPO, which samples $G$ complete trajectories and computes trajectory-level advantages $\hat{A}_{i}=R(\tau_{i})-\frac{1}{G}\sum_{l}R(\tau_{l})$; and (B) our truncated method, which fixes a prefix $\tau_{<t}$ and samples $k$ actions at step $t$ with step-level advantages $\hat{A}_{t}^{(j)}=r_{t}^{(j)}-\frac{1}{k}\sum_{l}r_{t}^{(l)}$ (see Appendix A.2 for the full gradient expressions).
Theorem 1 (Variance Reduction via Truncated Sampling).
Let $\tau=(a_{1},\ldots,a_{T})$ be a $T$-step trajectory. Suppose the trajectory-level reward decomposes additively as $R(\tau)=\sum_{t=1}^{T}r_{t}(a_{t},\tau_{<t})$, where $r_{t}$ is the step-$t$ reward. Assume the following conditions hold: 1) Non-negative future covariance: for each step $t$ and any fixed prefix $\tau_{<t}$, the covariance between the current step reward $r_{t}$ and the sum of future rewards $F_{t}$ satisfies $\text{Cov}(r_{t},F_{t}\mid\tau_{<t})\geq 0$. 2) Conditional independence: Step rewards are conditionally independent given the prefix trajectory. 3) Variance symmetry: Step rewards have comparable variance across steps, i.e., $\mathbb{E}_{\tau_{<t}}[\text{Var}[r_{t}\mid\tau_{<t}]]\approx\bar{v}$ for all $t$.
Then the per-sample variance of the scalar advantage in the truncated estimator satisfies:
$\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[\hat{A}_{t}^{(j)}\mid\tau_{<t}]\right]\leq\text{Var}\!\left[\hat{A}_{i}\right],$ (8)
where the left side is the expected (over prefixes) per-sample variance in the truncated estimator and the right side is the per-sample variance in the full-trajectory estimator. This holds under Assumption 1 alone. Moreover, under all three assumptions with equal group sizes $k=G$:
$\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[\hat{A}_{t}^{(j)}\mid\tau_{<t}]\right]\leq\frac{1}{T}\cdot\text{Var}\!\left[\hat{A}_{i}\right].$ (9)
Proof Sketch.
The proof proceeds in two parts (full details in Appendix A.3).
Part 1 (General bound, Eq. 8). Fixing the prefix $\tau_{<t}$ eliminates all randomness except the action at step $t$. By the law of total variance, the expected conditional variance $\mathbb{E}_{\tau_{<t}}[\text{Var}[R(\tau)\mid\tau_{<t}]]$ is at most $\text{Var}[R(\tau)]$. Because past rewards are constant given $\tau_{<t}$, the trajectory reward decomposes as $R(\tau)\mid\tau_{<t}=c+r_{t}+F_{t}$, and Assumption 1 ($\text{Cov}(r_{t},F_{t}\mid\tau_{<t})\geq 0$) ensures $\text{Var}[r_{t}\mid\tau_{<t}]\leq\text{Var}[R(\tau)\mid\tau_{<t}]$. Combining these bounds with $k=G$ yields Eq. 8.
Part 2 ($T$-fold reduction, Eq. 9). Under conditional independence (Assumption 2), the trajectory variance decomposes as $\text{Var}[R(\tau)]\geq\sum_{t}\mathbb{E}_{\tau_{<t}}[\text{Var}[r_{t}\mid\tau_{<t}]]$. Variance symmetry (Assumption 3) then gives $\mathbb{E}_{\tau_{<t}}[\text{Var}[r_{t}\mid\tau_{<t}]]\leq\frac{1}{T}\text{Var}[R(\tau)]$, which combined with Part 1 yields the $1/T$ factor. As a corollary, we show that truncated sampling also yields a $T$-fold reduction in total token generation cost to achieve the same advantage variance as standard GRPO (Proposition 2 in Appendix A.4). ∎
Since the policy gradient is a linear function of these advantages, lower advantage variance directly translates to lower-variance gradient estimates, enabling faster convergence and better final solutions (see Appendix A.6 for a detailed discussion).
Remark 1 (Credit Assignment).
The reward design provides an orthogonal benefit. By the data processing inequality, a binary EM reward provides at most 1 bit about the joint outcome of all $T$ actions, severely diluting the signal for any individual step. In contrast, the step-level LLM-judge reward $r_{t}$ directly evaluates $a_{t}$ in context, providing a substantially richer signal. Our ablation (Table 2) confirms that removing dense rewards causes larger drops than removing truncated sampling. We discuss the bias-variance trade-off of LLM-judge rewards in Appendix A.7.
5 Experiments
5.1 Datasets
We evaluate Slate on seven benchmark datasets spanning two categories: (1) General Question Answering: NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2022). (2) Multi-Hop Question Answering: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), Musique (Trivedi et al., 2022b), and Bamboogle (Press et al., 2022). These datasets encompass a diverse range of search-with-reasoning challenges, enabling comprehensive evaluation across both single-turn and multi-hop retrieval scenarios.
5.2 Baselines
We compare Slate against the following baselines: Inference without Retrieval: Direct generation and Chain-of-Thought (CoT) reasoning (Wei et al., 2022). Inference with Retrieval: RAG (Lewis et al., 2020), IRCoT (Trivedi et al., 2022a), and Search-o1 (Li et al., 2025). Fine-Tuning Methods: Supervised fine-tuning (SFT) (Chung et al., 2024) and RL without search (R1) (Guo et al., 2025). Search with RL: Search-R1 (Jin et al., 2025) (sparse outcome reward), ZeroSearch (Sun et al., 2025), ReSearch (Chen et al., 2025), and StepSearch (Wang et al., 2025) (step-wise process rewards).
5.3 Experimental Setup
We conduct experiments using Qwen2.5-7B-Base and Qwen2.5-3B-Base (Yang et al., 2024). For retrieval, we use the 2018 Wikipedia dump (Karpukhin et al., 2020) as the knowledge source and E5 (Wang et al., 2022) as the retriever, retrieving the top 3 passages per query. Following Search-R1, we merge the training sets of NQ and HotpotQA to form the unified training dataset. Exact Match (EM) is used as the primary evaluation metric (Yu et al., 2024). We use GRPO as the base RL algorithm with a policy learning rate of $1\times 10^{-6}$, $k=5$ truncated samples per step, clip ratio $\epsilon=0.2$, KL coefficient $\beta=0.001$, and early-termination bonus $\lambda=0.1$. For trajectory construction we use reward-weighted sampling (Section 3.2) with temperature $\eta=0.7$ to select which action extends the prefix at each step. The LLM-as-judge reward model uses Gemma3-27B as the evaluator. Training is performed for 500 steps on two NVIDIA A100 GPUs using LoRA (Hu et al., 2022) (rank 16, $\alpha=64$) for parameter-efficient fine-tuning in bfloat16 precision, with a batch size of 32, maximum sequence length of 4096 tokens, and maximum action budget $B=4$. Retrieved token loss masking is applied following Search-R1.
5.4 Main Results
The main results comparing Slate with all baselines across seven datasets are presented in Tables 1 and 3. We make the following key observations:
(1) Slate consistently achieves the best performance across all benchmarks. On the 7B model, Slate obtains an average EM of 0.461, representing a 3.0% absolute (7.0% relative) improvement over Search-R1 (0.431) and outperforming the best prior results on every individual dataset. On the 3B model, the improvement over Search-R1 is even more substantial (0.396 vs. 0.303, a 30.7% relative improvement), demonstrating that smaller models benefit more from the dense step-level supervision.
(2) Improvements are largest on hard multi-hop benchmarks. The gains of Slate over prior methods are non-uniform and scale with task difficulty. On the harder out-of-domain multi-hop datasets, Slate achieves the largest absolute improvements: on Musique (7B), the gain over Search-R1 is +5.1% and over StepSearch +3.1%; on Bamboogle (7B), +6.2% and +2.7%, respectively. On 2WikiMultiHopQA (7B), Slate obtains 0.413 vs. StepSearch’s 0.385 (+2.8%) and Search-R1’s 0.382 (+3.1%). This pattern is consistent with our hypothesis that dense step-level rewards help most when complex multi-step reasoning is required, as the credit assignment problem is most severe for longer trajectories. Notably, Slate is the only method that consistently outperforms both Search-R1 and StepSearch across all four multi-hop benchmarks, as prior methods show complementary strengths (e.g., Search-R1 excels on HotpotQA while StepSearch excels on 2Wiki and Bamboogle).
(3) Modest but consistent gains on general QA. On the general QA benchmarks, Slate outperforms Search-R1 by 1.3–1.9% absolute EM, with smaller gains expected for shorter trajectories where credit assignment is less challenging.
(4) Smaller models benefit proportionally more from dense supervision. On the 3B model (Table 3), gains over Search-R1 are dramatically larger on multi-hop benchmarks (e.g., +16.7% on Musique, +27.3% on Bamboogle), suggesting smaller models benefit most from explicit step-level supervision.
Table 1: Main results (Exact Match) on Qwen2.5-7B-Base across seven QA benchmarks. Best results are bolded, second best are underlined. Search-R1 and Slate are trained on NQ+HotpotQA. StepSearch is trained on MuSiQue (19k) and does not report general QA results. Evaluated on Wiki-18 knowledge base.
Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
Direct 0.213 0.512 0.212 0.187 0.246 0.027 0.123 0.217
CoT 0.228 0.529 0.231 0.217 0.261 0.045 0.168 0.240
RAG 0.371 0.582 0.339 0.287 0.231 0.061 0.214 0.298
IRCoT 0.342 0.561 0.312 0.141 0.142 0.072 0.216 0.255
Search-o1 0.321 0.548 0.298 0.193 0.181 0.053 0.302 0.271
SFT 0.353 0.569 0.325 0.196 0.269 0.054 0.110 0.268
R1 (no search) 0.298 0.541 0.278 0.234 0.270 0.076 0.287 0.283
ZeroSearch 0.411 0.601 0.389 0.294 0.275 0.102 0.258 0.333
ReSearch 0.398 0.594 0.372 0.294 0.264 0.143 0.373 0.348
Search-R1 0.480 0.638 0.457 0.433 0.382 0.196 0.432 0.431
StepSearch 0.380 0.385 0.216 0.467
Slate (Ours) 0.497 0.652 0.470 0.451 0.413 0.247 0.494 0.461
5.5 Analysis
5.5.1 Ablation Study
Table 2: Ablation study on Qwen2.5-7B-Base (Exact Match).
Variant HotpotQA 2Wiki Musique Bamboogle Avg.
Search-R1 (baseline) 0.433 0.382 0.196 0.432 0.361
(a) w/o truncated sampling 0.443 0.401 0.236 0.481 0.390
(b) w/o LLM-judge rewards 0.440 0.393 0.218 0.457 0.377
(c) EM reward only (final step) 0.437 0.387 0.205 0.444 0.368
Slate (full) 0.451 0.413 0.247 0.494 0.401
To understand the contribution of each component, we conduct an ablation study on Qwen2.5-7B-Base (Table 2). We evaluate four variants: (a) Slate without truncated sampling (using $k$ full trajectories with LLM-judge rewards), (b) Slate without LLM-judge rewards (using truncated sampling with sparse EM reward), (c) Slate with EM reward only at the final step (no thinking/query rewards), and (d) the full Slate system. The results reveal that both components contribute meaningfully, with their benefits most pronounced on harder datasets. Removing truncated sampling (variant a) reduces average EM by 1.1%, with the largest drops on the hardest benchmarks (Musique -1.1%, Bamboogle -1.3%), confirming that step-level variance reduction from shared prefixes improves optimization quality especially for complex multi-hop reasoning. Removing LLM-judge rewards (variant b) causes a larger drop of 2.4% average EM, demonstrating that dense step-level rewards provide a stronger training signal than sparse EM. The impact is again largest on the harder datasets (Musique -2.9%, Bamboogle -3.7%). Variant (c), which uses truncated sampling but only provides EM reward at the final step, performs close to Search-R1 (0.368 vs. 0.361 avg.), showing that truncated sampling alone without dense rewards provides limited benefit, it is the combination of both that yields the full improvement.
Figure 2: Training dynamics comparison on Qwen2.5-7B-Base. Slate converges faster and achieves a higher, more stable reward compared to Search-R1/GRPO and StepSearch/StePPO.
5.5.2 Training Dynamics
We compare the training reward curves of Slate against Search-R1 (GRPO) and StepSearch (StePPO) in Figure 2. Slate exhibits three notable properties: (1) Faster convergence: Slate reaches its peak training reward approximately 20% faster than StepSearch, attributable to the denser gradient signal from step-level rewards. (2) Higher reward ceiling: The final training reward of Slate is consistently higher than both baselines, reflecting the improved credit assignment from truncated sampling. (3) Greater stability: Unlike GRPO which can exhibit reward collapse, Slate maintains stable optimization throughout training due to the lower-variance advantage estimates predicted by Theorem 1.
5.5.3 Effect of Group Size $k$
We study the impact of $k\in\{1,3,5,7\}$ truncated samples per step (Appendix A.9). Performance improves steadily from k=1k{=}1 to k=5k{=}5 with diminishing returns at k=7k{=}7, consistent with the $1/k$ variance reduction from Eq. 4.
6 Conclusion
We presented Slate, a novel training method for search-augmented reasoning that introduces truncated step-level sampling and dense LLM-as-judge rewards. By sampling $k$ trajectory continuations that share a common prefix and differ only at the current decision step, our method achieves provably lower advantage variance than standard full-trajectory sampling, with up to a $T$-fold reduction for $T$-step trajectories, leading to lower-variance policy gradients. The dense LLM-as-judge rewards replace sparse binary outcome signals with rich, fine-grained step-level supervision that directly evaluates reasoning quality, query quality, and answer correctness on a ternary scale.
Our theoretical analysis provides formal guarantees on variance reduction, and our empirical results demonstrate significant improvements over both sparse-reward methods (Search-R1) and process-reward methods (StepSearch) across seven question answering benchmarks. Slate establishes a new state of the art for RL-based search-augmented reasoning, demonstrating that the combination of step-level exploration and dense model-based rewards is a powerful paradigm for training LLMs that reason effectively with external knowledge sources.
Appendix A Appendix
A.1 Algorithm
Algorithm 1 Step-Level Sampling with Dense LLM-Judge Rewards
1:Policy $\pi_{\theta}$, search engine $\mathcal{E}$, LLM judge $\mathcal{R}$, dataset $\mathcal{D}$, group size $k$, temperature $\eta$, max steps $B$, termination bonus $\lambda$
2:for each question $x\sim\mathcal{D}$ do
[…]
4:while $t\leq B$ do
5:  for $j=1,\ldots,k$ do $\triangleright$ Step-level sampling
6:   Sample $a_{t}^{(j)}=(s_{t}^{(j)},q_{t}^{(j)})\sim\pi_{\theta}(\cdot\mid x,\tau_{<t})$
7:   $r_{\text{think}}^{(j)}\leftarrow\mathcal{R}^{\text{think}}(s_{t}^{(j)},\tau_{<t})$
8:   if $a_{t}^{(j)}$ contains <search> then
9:    $r_{t}^{(j)}\leftarrow r_{\text{think}}^{(j)}+\mathcal{R}^{\text{query}}(q_{t}^{(j)},s_{t}^{(j)},\tau_{<t})$
10:   else if $a_{t}^{(j)}$ contains <answer> then
11:    $r_{t}^{(j)}\leftarrow r_{\text{think}}^{(j)}+\mathcal{R}^{\text{ans}}(a^{(j)},a_{\text{gold}},\tau_{<t})+\lambda\cdot\frac{B-t}{B}$
12:   end if
13:  end for
[…]
15:  $\bar{r}_{t}\leftarrow\frac{1}{k}\sum_{j}r_{t}^{(j)}$,   $\sigma_{t}\leftarrow\text{std}(\{r_{t}^{(j)}\})$
16:  for $j=1,\ldots,k$ do
17:   $\hat{A}_{t}^{(j)}\leftarrow(r_{t}^{(j)}-\bar{r}_{t})/(\sigma_{t}+\epsilon)$
18:  end for
[…]
20:  $\theta\leftarrow\theta+\alpha\nabla_{\theta}\Big[\frac{1}{k}\sum_{j=1}^{k}\mathcal{J}_{t}^{(j)}(\theta)-\beta\,\mathbb{D}_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\Big]$
21:   $\mathcal{J}_{t}^{(j)}(\theta)\!=\!\frac{1}{\sum_{l}I(y_{l})}\!\!\sum_{l:\,I(y_{l})=1}\!\!\min\!\big(\rho_{l}\hat{A}_{t}^{(j)},\;\text{clip}(\rho_{l},1{-}\epsilon,1{+}\epsilon)\hat{A}_{t}^{(j)}\big)$
22:   $\rho_{l}=\pi_{\theta}(y_{l}\mid y_{<l})/\pi_{\theta_{\text{old}}}(y_{l}\mid y_{<l})$,   $I(y_{l})=\mathbbm{1}[y_{l}\text{ is LLM-generated}]$
[…]
24:  $j^{*}\sim\text{softmax}\!\big(\hat{A}_{t}^{(1)},\ldots,\hat{A}_{t}^{(k)}\,/\,\eta\big)$
25:  $\tau_{<t+1}\leftarrow\tau_{<t}\cup\{a_{t}^{(j^{*})},\mathcal{E}(q_{t}^{(j^{*})})\}$
26:  $t\leftarrow t+1$
27:  if $a_{t}^{(j^{*})}$ contains <answer> then
28:   break
29:  end if
30:end while
31:end for
Table 3: Main results (Exact Match) on Qwen2.5-3B-Base across seven QA benchmarks. Best results are bolded, second best are underlined. Search-R1 and Slate are trained on NQ+HotpotQA. StepSearch is trained on MuSiQue (19k) and does not report general QA results. Evaluated on Wiki-18 knowledge base.
Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
Direct 0.152 0.434 0.178 0.145 0.249 0.018 0.030 0.172
CoT 0.168 0.451 0.196 0.163 0.257 0.032 0.072 0.191
RAG 0.312 0.504 0.296 0.251 0.221 0.051 0.076 0.244
IRCoT 0.286 0.478 0.268 0.171 0.196 0.059 0.245 0.243
Search-o1 0.267 0.465 0.254 0.240 0.207 0.045 0.316 0.256
SFT 0.301 0.491 0.286 0.191 0.248 0.039 0.112 0.238
R1 (no search) 0.258 0.467 0.243 0.191 0.277 0.051 0.232 0.246
ZeroSearch 0.348 0.531 0.365 0.260 0.234 0.056 0.096 0.270
ReSearch 0.339 0.518 0.342 0.261 0.228 0.074 0.184 0.278
Search-R1 0.406 0.587 0.435 0.284 0.273 0.049 0.088 0.303
StepSearch 0.329 0.339 0.181 0.328
Slate (Ours) 0.425 0.603 0.451 0.351 0.368 0.216 0.361 0.396
A.2 Gradient Estimators
The two gradient estimation strategies compared in Section 4 are defined as follows. For full-trajectory sampling (GRPO), the gradient estimate for step $t$ in trajectory $i$ is:
$\hat{g}_{t}^{\text{GRPO}}=\frac{1}{G}\sum_{i=1}^{G}\hat{A}_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(a_{t,i}\mid\tau_{<t,i}),$ (10)
where $\hat{A}_{i}=R(\tau_{i})-\frac{1}{G}\sum_{l}R(\tau_{l})$. For truncated step-level sampling (Slate), the gradient at step $t$ is:
$\hat{g}_{t}^{\text{Ours}}=\frac{1}{k}\sum_{j=1}^{k}\hat{A}_{t}^{(j)}\cdot\nabla_{\theta}\log\pi_{\theta}(a_{t}^{(j)}\mid\tau_{<t}),$ (11)
where $\hat{A}_{t}^{(j)}=r_{t}^{(j)}-\frac{1}{k}\sum_{l}r_{t}^{(l)}$.
A.3 Proof of Theorem 1
Proof.
We decompose the proof into two parts. Part 1 shows that truncated sampling never increases variance in expectation over prefixes (using Assumption 1), and Part 2 quantifies the improvement as a $T$-fold reduction under all three assumptions.
Part 1: General Variance Bound.
The core intuition is that fixing the prefix $\tau_{<t}$ eliminates all sources of randomness except the action at step $t$, reducing the variance of the advantage estimate on average.
Consider the full-trajectory advantage $\hat{A}_{i}=R(\tau_{i})-\bar{R}$, where $\bar{R}=\frac{1}{G}\sum_{i}R(\tau_{i})$. The variance of $\hat{A}_{i}$ taken over random trajectories $\tau_{i}\sim\pi_{\theta}(\cdot\mid x;\mathcal{E})$ depends on the total variability of $R(\tau)$:
$\text{Var}[\hat{A}_{i}]$ $=\text{Var}[R(\tau_{i})-\bar{R}]=\left(1-\tfrac{1}{G}\right)\text{Var}[R(\tau)].$
By the law of total variance applied to the trajectory prefix $\tau_{<t}$:
$\text{Var}[R(\tau)]$ $=\underbrace{\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right]}_{\text{within-prefix variance}}\quad+\underbrace{\text{Var}_{\tau_{<t}}\!\left[\mathbb{E}[R(\tau)\mid\tau_{<t}]\right]}_{\text{between-prefix variance}\;\geq\;0}.$
Since both terms on the right are non-negative, the expected conditional variance is bounded by the unconditional variance:
$\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right]\leq\text{Var}[R(\tau)].$ (12)
Now, in our method, the step-level advantage $\hat{A}_{t}^{(j)}$ is computed with the prefix $\tau_{<t}$ fixed. Its conditional variance is:
$\text{Var}[\hat{A}_{t}^{(j)}\mid\tau_{<t}]=\left(1-\frac{1}{k}\right)\text{Var}[r_{t}\mid\tau_{<t}].$ (13)
We now show that $\text{Var}[r_{t}\mid\tau_{<t}]\leq\text{Var}[R(\tau)\mid\tau_{<t}]$ for each prefix. Given a fixed prefix $\tau_{<t}$, the rewards from steps $1,\ldots,t{-}1$ are constants, so $R(\tau)\mid\tau_{<t}=c+r_{t}+F_{t}$ where $c=\sum_{t^{\prime}<t}r_{t^{\prime}}$ is constant and $F_{t}=\sum_{t^{\prime}=t+1}^{T}r_{t^{\prime}}$ denotes the future rewards. Therefore:
$\text{Var}[R(\tau)\mid\tau_{<t}]$ $=\text{Var}[r_{t}+F_{t}\mid\tau_{<t}]$
$=\text{Var}[r_{t}\mid\tau_{<t}]+\text{Var}[F_{t}\mid\tau_{<t}]+2\text{Cov}(r_{t},F_{t}\mid\tau_{<t}).$ (14)
Under Assumption 1 ($\text{Cov}(r_{t},F_{t}\mid\tau_{<t})\geq 0$) and since $\text{Var}[F_{t}\mid\tau_{<t}]\geq 0$, we obtain:
$\text{Var}[r_{t}\mid\tau_{<t}]\leq\text{Var}[R(\tau)\mid\tau_{<t}].$ (15)
Taking expectations over prefixes and applying Eq. 12:
$\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right]\leq\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right]\leq\text{Var}[R(\tau)].$ (16)
Combining these and setting $k=G$:
$\mathbb{E}_{\tau_{<t}}\!\big[\text{Var}[\hat{A}_{t}^{(j)}\mid\tau_{<t}]\big]$ $=\left(1-\tfrac{1}{k}\right)\mathbb{E}_{\tau_{<t}}\!\big[\text{Var}[r_{t}\mid\tau_{<t}]\big]$
$\leq\left(1-\tfrac{1}{G}\right)\text{Var}[R(\tau)]=\text{Var}[\hat{A}_{i}].$ (17)
This establishes that, with equal group sizes, the truncated estimator has no more variance in expectation than the full-trajectory estimator.
Part 2: $T$-fold Reduction Under Independence and Symmetry.
Part 1 shows the truncated estimator is never worse; we now show it can be $T$ times better. The intuition is that the total trajectory variance $\text{Var}[R(\tau)]$ is the sum of variances from all $T$ steps, but the truncated estimator only “sees” the variance from one step—hence the $1/T$ factor.
Under Assumption 2 (conditional independence), for each step $t$, the reward $r_{t}$ depends only on $a_{t}$ and $\tau_{<t}$, and future actions do not affect past rewards. Under this condition, the trajectory reward variance decomposes as:
$\text{Var}[R(\tau)]$ $=\text{Var}\!\left[\sum_{t=1}^{T}r_{t}\right]=\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right]+\text{inter-step terms}.$
Under Assumption 2, the inter-step covariance terms are non-negative: conditional independence ensures that $\text{Cov}(r_{t},r_{t^{\prime}}\mid\tau_{<t})\geq 0$ for $t<t^{\prime}$, since any remaining dependence flows only through the sequential prefix structure. Therefore:
$\text{Var}[R(\tau)]\geq\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right].$ (18)
Under Assumption 3 (variance symmetry), $\mathbb{E}_{\tau_{<t}}[\text{Var}[r_{t}\mid\tau_{<t}]]\approx\bar{v}$ for all $t$, so the right-hand side simplifies to $T\cdot\bar{v}$, giving:
$\text{Var}[R(\tau)]\geq T\cdot\bar{v}\geq T\cdot\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right].$ (19)
Rearranging yields $\mathbb{E}_{\tau_{<t}}[\text{Var}[r_{t}\mid\tau_{<t}]]\leq\frac{1}{T}\text{Var}[R(\tau)]$. Combining with the result from Part 1 (with $k=G$):
$\mathbb{E}_{\tau_{<t}}\!\big[\text{Var}[\hat{A}_{t}^{(j)}\mid\tau_{<t}]\big]$ $=\left(1-\tfrac{1}{G}\right)\mathbb{E}_{\tau_{<t}}\!\big[\text{Var}[r_{t}\mid\tau_{<t}]\big]$
$\leq\left(1-\tfrac{1}{G}\right)\frac{1}{T}\text{Var}[R(\tau)]$
$=\frac{1}{T}\cdot\text{Var}[\hat{A}_{i}].$ (20)
A.4 Sample Efficiency
Proposition 2 (Sample Efficiency).
To achieve the same advantage variance as standard GRPO with $G$ full-trajectory samples, the truncated step-level method requires only $G/T$ samples per step under the conditions of Theorem 1, yielding a $T$-fold reduction in total token generation cost.
A.5 Proof of Proposition 2
Proof.
The variance of each method’s advantage estimator for step $t$ scales inversely with the number of samples. For standard GRPO, the variance of the mean advantage estimator is $\frac{1}{G}\text{Var}[\hat{A}_{i}]$. For the truncated method, the expected conditional variance of the mean advantage estimator is:
$\frac{1}{k}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[\hat{A}_{t}^{(j)}\mid\tau_{<t}]\right]\leq\frac{1}{k}\cdot\frac{1}{T}\cdot\text{Var}[\hat{A}_{i}],$ (21)
where the inequality follows from Theorem 1. Equating the two to find the minimum $k$:
$\frac{1}{k}\cdot\frac{1}{T}\cdot\text{Var}[\hat{A}_{i}]=\frac{1}{G}\cdot\text{Var}[\hat{A}_{i}]\quad\Longrightarrow\quad k=\frac{G}{T}.$ (22)
Thus, only $G/T$ samples per step suffice.
We now count total tokens generated. In standard GRPO, we sample $G$ complete trajectories of average length $L$, costing $G\cdot L$ tokens. In our method, each truncated sample generates only one step’s worth of tokens—approximately $L/T$ tokens, since the full trajectory’s $L$ tokens are spread over $T$ steps—and we repeat this sampling at each of the $T$ steps. The total cost is therefore:
[…]
A.6 Variance Reduction and Convergence
Theorem 1 establishes that the truncated estimator yields lower-variance scalar advantages $\hat{A}_{t}^{(j)}$. Since the policy gradient $\hat{g}_{t}=\frac{1}{k}\sum_{j}\hat{A}_{t}^{(j)}\nabla_{\theta}\log\pi_{\theta}(a_{t}^{(j)}\mid\tau_{<t})$ is a linear function of these advantages, lower advantage variance directly translates to lower variance in the gradient estimates (modulated by the score function magnitudes).
Intuitively, the policy gradient estimate $\hat{g}$ acts as a noisy compass: it points toward the true gradient on average, but individual estimates may deviate substantially. In standard GRPO, the gradient signal for step $t$ is weighted by the trajectory-level advantage $\hat{A}_{i}$, which conflates the quality of all $T$ actions. A trajectory may succeed despite a poor intermediate step, causing that step to be incorrectly reinforced; conversely, a trajectory may fail despite strong intermediate reasoning, penalizing all steps indiscriminately. These “mislabeled” updates are the practical manifestation of high advantage variance. Over many updates they cancel in expectation, but each wasted update consumes compute budget and slows progress. High variance also forces the use of smaller learning rates to avoid divergence, further limiting the speed of convergence. In our truncated method, fixing the prefix $\tau_{<t}$ and varying only step $t$ ensures that the advantage $\hat{A}_{t}^{(j)}$ reflects solely the quality of the current action, so every gradient update sends an accurate signal. Beyond faster convergence, lower variance can also lead to better final solutions: cleaner gradients allow the optimizer to reliably descend into sharper, higher-performing regions of the loss landscape that noisy updates would overshoot or bounce out of, and they ensure that more of the finite training budget contributes useful learning signal rather than noise.
A.7 Bias-Variance Trade-off
Remark 2 (Bias-Variance Trade-off).
Our step-level LLM-judge rewards may introduce bias if the judge does not perfectly predict the contribution of step $t$ to the final outcome. However, this bias is typically small for well-calibrated LLM judges, and the significant variance reduction (Theorem 1) more than compensates, leading to faster convergence and better final performance. This aligns with the classical bias-variance trade-off in policy gradient methods, where moderate bias with substantially reduced variance yields better optimization dynamics (Schulman et al., 2015).
A.8 Early-Termination Bonus Details
The bonus term $\lambda\cdot(B-t)/B$ in Eq. 6 encourages the model to produce an answer as soon as it has gathered sufficient information. Without such a term, the model may learn to issue superfluous search queries—each receiving a neutral or mildly positive reward from the LLM judge—even when the information needed to answer has already been retrieved. The bonus is largest when the model answers early (e.g., $\lambda\cdot\frac{3}{4}$ at step t=1t{=}1 with B=4B{=}4) and zero when it exhausts the full budget (t=Bt{=}B), creating a progressive incentive to terminate sooner. Crucially, this bonus only applies to answer candidates, so at any step where some of the $k$ sampled actions produce an answer and others produce a search query, the answer candidates receive a higher reward, creating a meaningful advantage signal that survives the group normalization in Eq. 4. This ensures the policy gradient directly reinforces early termination when additional search steps are unlikely to improve the answer.
A.9 Effect of Group Size $k$
We study the impact of the number of truncated samples $k\in\{1,3,5,7\}$ per step on Qwen2.5-7B-Base (Table 4). When $k=1$, the method reduces to standard REINFORCE with LLM-judge rewards (no group-relative advantage). Performance improves steadily from $k=1$ to $k=5$, with diminishing returns at $k=7$. This is consistent with our theoretical analysis: increasing $k$ reduces the variance of the step-level advantage estimate (Eq. 4), but the marginal benefit decreases as $1/k$.
Table 4: Effect of group size $k$ (truncated samples per step) on Qwen2.5-7B-Base (EM).
Group Size $k$ HotpotQA 2Wiki Musique Bamboogle Avg.
$k=1$ 0.437 0.390 0.213 0.452 0.373
$k=3$ 0.446 0.405 0.237 0.483 0.393
$k=5$ (default) 0.451 0.413 0.247 0.494 0.401
$k=7$ 0.449 0.411 0.244 0.491 0.399
A.10 LLM-as-Judge Reward Prompts
We provide the exact prompts used for the three LLM-as-judge reward components described in Section 3.3. In each prompt, the placeholders in curly braces (e.g., {context}, {thinking}) are filled with the corresponding trajectory content at evaluation time. The judge is instructed to produce a chain-of-thought explanation before the score, enclosed in XML-style tags.
Thinking Reward Prompt.
Evaluate the quality of the following reasoning step in a search-based question answering system.
Context: {context}
Current Thinking Step: {thinking}
The reasoning should be based on the previous context and the question, nothing else.
Evaluate this thinking step on these criteria:
1. Relevance: Does it address the question appropriately?
2. Clarity: Is the reasoning clear and logical?
3. Specificity: Does it identify concrete information needs?
4. Progress: Does it move toward answering the question?
5. Faithfulness: Does it accurately reflect the information in the previous context? Is there any out-of-context information?
Provide a score using EXACTLY one of these three values:
- +1: GOOD -- Clear, relevant reasoning that identifies specific information needs and moves toward answering the question
-  0: ACCEPTABLE -- Reasoning is somewhat relevant but vague, lacks specificity, or makes only minimal progress
- -1: BAD -- Irrelevant, misleading, or counterproductive reasoning that does not help answer the question
First provide your reasoning, then the score. Use this exact format:
<explanation> Your reasoning here </explanation>
<score> numerical score </score>
Query Generation Reward Prompt.
Evaluate the quality of the following search query for a question answering system.
Context: {context}
Thinking before this query: {thinking}
Generated Query: {query}
IMPORTANT: This is a multi-step reasoning system. The query does NOT need to directly answer the final question in one step. Instead, evaluate whether it makes good progress toward the answer by retrieving useful intermediate information.
Evaluate this query on these criteria:
1. Relevance: Will it retrieve information that makes progress toward answering the question? (Intermediate steps are valuable!)
2. Specificity: Is it specific enough to get useful results?
3. Searchability: Is it well-formed for a search engine with appropriate keywords? Good queries combine multiple relevant terms.
4. Alignment: Does it align with the thinking step that preceded it?
5. Novelty: Does it explore new information (not redundant with the context)? If the context already contains the answer to what the query is searching for, the query is redundant and unhelpful.
Provide a score using EXACTLY one of these three values:
- +1: GOOD -- Specific, well-formed query that will retrieve useful information to make progress (even if intermediate). Has clear keywords and good searchability. Combines multiple relevant terms or uses specific names/concepts.
-  0: ACCEPTABLE -- Query has some specificity but could be improved. May lack context-specific keywords or be somewhat generic, but shows reasonable attempt at targeting the information need.
- -1: BAD -- Single generic word without context (e.g., just ‘‘singer’’, ‘‘perfume’’, ‘‘city’’), completely irrelevant to the question, redundant with information already in the context, or so poorly formed it will return millions of unhelpful results.
First provide your reasoning, then the score. Use this exact format:
<explanation> Your reasoning here </explanation>
<score> numerical score </score>
Final Answer Reward Prompt.
Evaluate if the predicted answer correctly answers the question.
Context: {context}
Ground Truth Answer: {ground_truth}
Predicted Answer: {predicted_answer}
Compare the predicted answer to the ground truth. They don’t need to be word-for-word identical, but the predicted answer should convey the same core information.
Provide a score using EXACTLY one of these three values:
- +1: CORRECT -- The predicted answer conveys the same core information as the ground truth
-  0: PARTIALLY CORRECT -- The answer is incomplete, ambiguous, or contains minor inaccuracies
- -1: INCORRECT -- The answer is wrong or contradicts the ground truth
First provide your reasoning, then the score. Use this exact format:
<explanation> Your reasoning here </explanation>
<score> numerical score </score>

### v2

@@
-arXiv:2602.23440v1[cs.CL] 26 Feb 2026
+Truncated Step-Level Sampling with Process Rewards for Retrieval-Augmented Reasoning
 Abstract
@@
 Process-reward methods like StepSearch alleviate this by introducing step-level supervision, but rely on heuristic rewards such as TF-IDF overlap with gold documents, and still sample $k$ complete trajectories per example, retaining high gradient variance.
-We propose Slate, a framework built on two complementary ideas: (1) truncated step-level sampling, which generates $k$ trajectories that share a common prefix and differ only at the next step, and (2) dense LLM-as-judge rewards, which replace heuristic scoring with a capable LLM evaluator that assesses the quality of each reasoning step, search query, and answer, providing richer and more reliable supervision.
-We theoretically prove that under the same dense reward structure, truncated sampling reduces the variance of advantage estimates by up to a factor of $T$ compared to full-trajectory sampling for $T$-step trajectories, yielding lower-variance, better-targeted policy gradients.
+We propose Slate, a framework built on two complementary ideas: (1) truncated step-level sampling, which generates $k$ trajectories that share a common prefix and differ only at the next step, isolating variation to a single decision point; and (2) dense, decomposed LLM-as-judge rewards, which score each reasoning step, search query, and answer on a ternary scale with separate quality dimensions, providing richer supervision than binary outcome signals or undifferentiated step-level judgments.
+We theoretically prove that under the same dense reward structure, truncated sampling reduces the variance of advantage estimates by up to a factor of $T$ compared to full-trajectory sampling for $T$-step trajectories, yielding lower-variance and better-targeted policy gradients.
 Experiments on seven QA benchmarks confirm that Slate consistently outperforms both sparse-reward and process-reward baselines, with the largest gains on harder multi-hop tasks and smaller models.
 1 Introduction
-Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation (Hendrycks et al., 2020; Clark et al., 2018).
+Large language models have demonstrated remarkable capabilities in natural language understanding and generation (Hendrycks et al., 2020; Clark et al., 2018).
 Despite these achievements, LLMs often struggle with complex reasoning tasks (Wei et al., 2022) and lack access to up-to-date external knowledge (Jin et al., 2024).
 Integrating search engines into the LLM reasoning loop, where the model interleaves its own chain-of-thought reasoning with external retrieval calls, has emerged as a promising paradigm for knowledge-intensive question answering (Yao et al., 2023; Trivedi et al., 2022a).
-Reinforcement learning (RL) provides a natural framework for optimizing such search-augmented reasoning systems.
-Search-R1 (Jin et al., 2025) pioneered the use of RL (PPO and GRPO) to train LLMs that autonomously invoke search engines during multi-turn reasoning, using outcome-based exact-match rewards.
-While effective, this sparse reward design suffers from a fundamental limitation: the credit assignment problem.
-When the model receives a single binary signal after completing a multi-step trajectory, potentially involving several rounds of thinking, query generation, and information processing, it cannot determine which intermediate steps contributed positively or negatively to the outcome.
-Recent work has attempted to address this limitation through process-level supervision.
-StepSearch (Wang et al., 2025) augments PPO with step-wise rewards based on information gain and redundancy penalties, achieving notable improvements on multi-hop QA benchmarks.
-However, StepSearch still relies on sampling $k$ complete trajectories per training example, which is computationally expensive and retains variance from all steps in the gradient estimates for any individual step.
+Reinforcement learning provides a natural framework for optimizing such systems.
+Search-R1 (Jin et al., 2025) pioneered RL training for LLMs that invoke search engines during multi-turn reasoning, using outcome-based exact-match rewards.
+While effective, this sparse reward suffers from the credit assignment problem: a single binary signal after a multi-step trajectory cannot attribute success or failure to individual steps.
+Process-level supervision methods, StepSearch (Wang et al., 2025) with heuristic step rewards, and SWiRL (Goldie et al., 2025) with LLM-judge binary step rewards, improve over outcome-only supervision, but both still sample complete trajectories independently, meaning step-level advantages conflate variation from the current action with variation across different prefix histories, leaving the credit assignment and gradient variance problems only partially resolved.
 In this paper, we propose Slate111code available at https://github.com/algoprog/SLATE, Step-Level Advantage estimation for Truncated Exploration, which addresses these limitations through two complementary ideas:
@@
   • Dense LLM-as-Judge Rewards: We replace the sparse EM outcome reward with dense, step-level rewards produced by an LLM evaluator.
-The judge scores each reasoning step (thinking quality), each search query (query quality), and the final answer (correctness) on a discrete scale $\{-1,0,+1\}$, providing rich supervision at every decision point.
-We provide formal theoretical analysis showing that truncated step-level sampling yields advantage estimates with provably lower variance than standard full-trajectory sampling, with up to a $T$-fold reduction for $T$-step trajectories, resulting in lower-variance policy gradients.
-Combined with the dense reward signal from the LLM judge, this enables faster convergence, better credit assignment, and improved final performance.
+The judge scores each reasoning step (thinking quality), each search query (query quality), and the final answer (correctness) on a discrete scale, providing rich supervision at every decision point.
 Our contributions are the following:
   1. 1.
-    We propose truncated step-level sampling for GRPO, an exploration strategy that generates $k$ alternative continuations only at the current decision point rather than $k$ complete trajectories.
+    We propose truncated step-level sampling for GRPO, an exploration strategy that generates $k$ alternative continuations from a shared prefix at the current decision point rather than $k$ complete trajectories.
 We formally prove this achieves a $T$-fold advantage variance reduction over full-trajectory sampling, yielding lower-variance policy gradients.
+This principled exploration mechanism is, to our knowledge, the first to provide formal guarantees on variance reduction for step-level RL in search-augmented reasoning.
   2. 2.
-    We introduce an LLM-as-judge reward system that provides dense, discrete step-level supervision for reasoning quality, search query quality, and answer correctness, replacing the sparse outcome reward.
+    We introduce a decomposed LLM-as-judge reward system with separate ternary scores for reasoning quality, query quality, and answer correctness, along with an early-termination bonus.
+Unlike prior binary step-level judgments (Goldie et al., 2025), this structured design provides richer supervision that disentangles the distinct skills required for effective retrieval-augmented reasoning.
   3. 3.
-    We demonstrate that Slate significantly outperforms both existing sparse-reward methods (Search-R1) and process-reward methods (StepSearch) on seven QA benchmarks.
+    We demonstrate that Slate significantly outperforms both existing sparse-reward methods (Search-R1) and process-reward methods (StepSearch) on seven QA benchmarks, and that our ablation confirms truncated sampling provides gains beyond what step-level rewards alone achieve.
 2 Related Work
 LLMs and Retrieval
-Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) integrates external knowledge into LLM generation by retrieving relevant passages based on the input query.
-While effective for simple lookups, standard RAG struggles with multi-hop reasoning where iterative retrieval is needed (Yang et al., 2018; Trivedi et al., 2022a).
-Alternative approaches treat the search engine as a tool (Yao et al., 2023; Schick et al., 2023), enabling the LLM to decide when and what to search.
-However, prompting-based tool-use methods rely on in-context examples and fail to generalize robustly, while supervised fine-tuning approaches require expensive annotated trajectories (Schick et al., 2023; Asai et al., 2024).
+Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) integrates external knowledge into LLM generation but struggles with multi-hop reasoning requiring iterative retrieval (Yang et al., 2018; Trivedi et al., 2022a).
+Tool-use approaches (Yao et al., 2023; Schick et al., 2023) let the LLM decide when and what to search, but prompting-based methods fail to generalize robustly and supervised fine-tuning requires expensive annotated trajectories (Schick et al., 2023; Asai et al., 2024).
 RL for LLM Reasoning
-Reinforcement learning has emerged as a potent paradigm for enhancing LLM reasoning, as demonstrated by OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025).
-Policy gradient methods including PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) have been widely adopted for training reasoning models.
-GRPO is particularly attractive as it eliminates the need for a separate critic model by using group-relative advantages computed from multiple sampled responses.
-However, applying these methods to search-augmented scenarios introduces unique challenges around reward design, credit assignment, and integration of retrieved content.
+Reinforcement learning has become a key paradigm for LLM reasoning (Jaech et al., 2024; Guo et al., 2025), with policy gradient methods such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) widely adopted.
+GRPO is particularly attractive as it eliminates the need for a separate critic by using group-relative advantages.
+However, applying these methods to search-augmented scenarios introduces challenges around reward design, credit assignment, and integration of retrieved content.
 Search-Augmented RL
@@
 StepSearch (Wang et al., 2025) addresses the reward sparsity problem by introducing step-wise rewards with information gain and redundancy penalties, but still samples complete trajectories and relies on access to ground-truth intermediate documents for computing information gain.
-Our work differs from all prior methods by combining truncated step-level sampling with model-based dense rewards from an LLM judge (eliminating the need for ground-truth intermediate annotations).
+SWiRL (Goldie et al., 2025) explored using an LLM judge to provide step-level binary rewards for multi-step reasoning with tool use.
+However, SWiRL operates offline, generating complete trajectories, segmenting them into sub-trajectories, and training on this static corpus, so step-level advantages still conflate variation across different prefix histories with variation at the current step.
+It also uses undifferentiated binary step judgments and provides no formal variance analysis.
+Our work differs along three axes: (1) truncated step-level sampling generates $k$ continuations from an identical shared prefix, isolating variation to exactly one decision point with provable variance guarantees (Theorem 1); (2) online GRPO optimization avoids the distribution mismatch of offline approaches; and (3) our decomposed ternary reward design provides richer supervision than heuristic step rewards (Wang et al., 2025) or binary LLM judgments (Goldie et al., 2025), without requiring ground-truth intermediate annotations.
 Figure 1: Comparison of GRPO (with full trajectory sampling) and our truncated step-level sampling. By fixing the prefix $\tau_{<t}$, all variation in the sampled group is localized to step $t$.
@@
 Query Reward.
-The query reward $r_{\text{query}}(q_{t},s_{t},\tau_{<t})\in\{-1,0,+1\}$ evaluates the search query $q_{t}$ conditioned on the retrieved documents, the preceding reasoning and the trajectory context.
+The query reward $r_{\text{query}}(q_{t},s_{t},\tau_{<t})\in\{-1,0,+1\}$ evaluates the search query $q_{t}$ conditioned on the preceding reasoning and the trajectory context.
 The judge assesses five criteria: relevance, specificity, searchability, alignment, and novelty.
 A score of +1+1 indicates a specific, well-formed query with clear keywords; 0 indicates some specificity but generic phrasing; and 1-1 indicates an irrelevant or poorly formed query.
+By design, the query reward evaluates the formulation quality of the query before observing retrieval results; this decoupling ensures the reward reflects intrinsic query properties (keyword specificity, search-engine compatibility) rather than fluctuating with retrieval nondeterminism or index coverage, while the downstream effect of retrieval quality is captured through subsequent steps and the final answer reward.
 Final Answer Reward.
@@
 $\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[\hat{A}_{t}^{(j)}\mid\tau_{<t}]\right]\leq\frac{1}{T}\cdot\text{Var}\!\left[\hat{A}_{i}\right].$ (9)
+Note: The two estimators target different quantities, the trajectory-level advantage $\hat{A}_{i}$ estimates the deviation of the full return $R(\tau)$, while the step-level advantage $\hat{A}_{t}^{(j)}$ estimates the deviation of the single-step reward $r_{t}$.
+The comparison below therefore characterizes a bias-variance trade-off: the step-level estimator achieves lower variance at the cost of discarding future-reward information (see Remark 1 for when this substitution is justified).
 Proof Sketch.
@@
 Since the policy gradient is a linear function of these advantages, lower advantage variance directly translates to lower-variance gradient estimates, enabling faster convergence and better final solutions (see Appendix A.7 for a detailed discussion).
-Remark 1 (Credit Assignment).
-The reward design provides an orthogonal benefit.
-By the data processing inequality, a binary EM reward provides at most 1 bit about the joint outcome of all $T$ actions, severely diluting the signal for any individual step.
-In contrast, the step-level LLM-judge reward $r_{t}$ directly evaluates $a_{t}$ in context, providing a substantially richer signal.
-Our ablation (Table 2) confirms that removing dense rewards causes larger drops than removing truncated sampling.
-We discuss the bias-variance trade-off of LLM-judge rewards in Appendix A.7.
+We further discuss the bias-variance trade-off of estimator substitution and the credit assignment benefits of dense rewards in Appendix A.4.
 5 Experiments
-5.1 Datasets
+Datasets
 We evaluate Slate on seven benchmark datasets spanning two categories: (1) General Question Answering: NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2022).
@@
 These datasets encompass a diverse range of search-with-reasoning challenges, enabling comprehensive evaluation across both single-turn and multi-hop retrieval scenarios.
-5.2 Baselines
+Baselines
 We compare Slate against the following baselines: Inference without Retrieval: Direct generation and Chain-of-Thought (CoT) reasoning (Wei et al., 2022).
@@
 Search with RL: Search-R1 (Jin et al., 2025) (sparse outcome reward), ZeroSearch (Sun et al., 2025), ReSearch (Chen et al., 2025), and StepSearch (Wang et al., 2025) (step-wise process rewards).
-5.3 Experimental Setup
+Experimental Setup
 We conduct experiments using Qwen2.5-7B-Base and Qwen2.5-3B-Base (Yang et al., 2024).
@@
 The main results comparing Slate with all baselines across seven datasets are presented in Tables 1 and 3. We make the following key observations:
-(1) Slate consistently achieves the best performance across all benchmarks.
+(1) Slate consistently performs best across all benchmarks.
 On the 7B model, Slate obtains an average EM of 0.461, representing a 3.0% absolute (7.0% relative) improvement over Search-R1 (0.431) and outperforming the best prior results on every individual dataset.
@@
 Notably, Slate is the only method that consistently outperforms both Search-R1 and StepSearch across all four multi-hop benchmarks, as prior methods show complementary strengths (e.g., Search-R1 excels on HotpotQA while StepSearch excels on 2Wiki and Bamboogle).
-(3) Modest but consistent gains on general QA.
-On the general QA benchmarks, Slate outperforms Search-R1 by 1.3–1.9% absolute EM, with smaller gains expected for shorter trajectories where credit assignment is less challenging.
-(4) Smaller models benefit proportionally more from dense supervision.
+(3) Gains on general QA and out-of-domain generalization.
+On the general QA benchmarks, Slate outperforms Search-R1 by 1.3–1.9% absolute EM.
+Notably, five of our seven evaluation benchmarks are out-of-domain (trained only on NQ+HotpotQA), and the largest gains occur on these unseen multi-hop tasks, demonstrating that step-level supervision teaches transferable reasoning skills rather than dataset-specific shortcuts.
+(4) Smaller models benefit more from dense supervision.
 On the 3B model (Table 3), gains over Search-R1 are dramatically larger on multi-hop benchmarks (e.g., +16.7% on Musique, +27.3% on Bamboogle), suggesting smaller models benefit most from explicit step-level supervision.
@@
 Slate (Ours) 0.497 0.652 0.470 0.451 0.413 0.247 0.494 0.461
-5.5 Analysis
-5.5.1 Ablation Study
+Ablation Study
 Table 2: Ablation study on Qwen2.5-7B-Base (Exact Match).
+Variant (a) mirrors the SWiRL (Goldie et al., 2025) approach of applying step-level LLM-judge rewards with full-trajectory sampling (adapted to our online setting).
 Variant HotpotQA 2Wiki Musique Bamboogle Avg.
@@
 Slate (full) 0.451 0.413 0.247 0.494 0.401
-To understand the contribution of each component, we conduct an ablation study on Qwen2.5-7B-Base (Table 2).
-We evaluate four variants: (a) Slate without truncated sampling (using $k$ full trajectories with LLM-judge rewards), (b) Slate without LLM-judge rewards (using truncated sampling with sparse EM reward), (c) Slate with EM reward only at the final step (no thinking/query rewards), and (d) the full Slate system.
-The results reveal that both components contribute meaningfully, with their benefits most pronounced on harder datasets.
-Removing truncated sampling (variant a) reduces average EM by 1.1%, with the largest drops on the hardest benchmarks (Musique -1.1%, Bamboogle -1.3%), confirming that step-level variance reduction from shared prefixes improves optimization quality especially for complex multi-hop reasoning.
-Removing LLM-judge rewards (variant b) causes a larger drop of 2.4% average EM, demonstrating that dense step-level rewards provide a stronger training signal than sparse EM.
-The impact is again largest on the harder datasets (Musique -2.9%, Bamboogle -3.7%).
-Variant (c), which uses truncated sampling but only provides EM reward at the final step, performs close to Search-R1 (0.368 vs.
-0.361 avg.), showing that truncated sampling alone without dense rewards provides limited benefit, it is the combination of both that yields the full improvement.
-Figure 2: Training dynamics comparison on Qwen2.5-7B-Base. Slate converges faster and achieves a higher, more stable reward compared to Search-R1/GRPO and StepSearch/StePPO.
-5.5.2 Training Dynamics
+Table 2 ablates each component of Slate on Qwen2.5-7B-Base.
+Variant (a) is particularly informative: it applies dense LLM-judge rewards with full-trajectory sampling, closely mirroring the SWiRL (Goldie et al., 2025) approach (adapted to our online setting).
+While this already improves over Search-R1 (+2.9% avg.), the full Slate with truncated sampling achieves a further +1.1% gain, with the largest improvements on the hardest benchmarks (Musique ++1.1%, Bamboogle ++1.3%).
+This demonstrates that truncated sampling provides gains above and beyond step-level rewards alone, confirming that the exploration strategy, not just the reward signal, is critical, consistent with Theorem 1.
+Removing LLM-judge rewards (variant b) causes a larger drop of 2.4% average EM, with the biggest impact on harder datasets (Musique -2.9%, Bamboogle -3.7%).
+Variant (c), with only EM reward at the final step, performs close to Search-R1 (0.368 vs.
+0.361).
+Note that variant (c) still receives a non-trivial training signal at intermediate steps: among the $k$ candidates sampled at each step, some may produce an <answer> (receiving the EM reward) while others produce a <search> (receiving zero reward), creating a nonzero advantage that trains the model on when to stop searching.
+However, the small margin over Search-R1 confirms that it is the synergy of both components that yields the full improvement.
+Figure 2: Training dynamics comparison on Qwen2.5-7B-Base.
+Training Dynamics
 We compare the training reward curves of Slate against Search-R1 (GRPO) and StepSearch (StePPO) in Figure 2.
@@
 (3) Greater stability: Unlike GRPO which can exhibit reward collapse, Slate maintains stable optimization throughout training due to the lower-variance advantage estimates predicted by Theorem 1.
-5.5.3 Effect of Group Size $k$
-We study the impact of $k\in\{1,3,5,7\}$ truncated samples per step (Appendix A.9).
-Performance improves steadily from k=1k{=}1 to k=5k{=}5 with diminishing returns at k=7k{=}7, consistent with the $1/k$ variance reduction from Eq.
-4.
+We also study the effect of the number of truncated samples $k$ per step in Appendix A.12.
+Performance improves steadily from $k=1$ to $k=5$ with diminishing returns at $k=7$, consistent with the $1/k$ variance reduction predicted by Theorem 1.
+Why Process Rewards Succeed in Search.
+In math, step-level rewards as RL training signals generally underperform outcome-based rewards (Uesato et al., 2022; Lightman et al., 2023).
+Search-augmented reasoning differs in several key ways: steps are externally grounded by retrieval results (eliminating capability mismatch), rewards are evaluative rather than predictive, and trajectories are short ($T\leq 4$) and modular.
+Truncated sampling further addresses prefix confounding by isolating variation to a single action (Theorem 1).
+We provide a detailed discussion in Appendix A.11.
 6 Conclusion
-We presented Slate, a novel training method for search-augmented reasoning that introduces truncated step-level sampling and dense LLM-as-judge rewards.
-By sampling $k$ trajectory continuations that share a common prefix and differ only at the current decision step, our method achieves provably lower advantage variance than standard full-trajectory sampling, with up to a $T$-fold reduction for $T$-step trajectories, leading to lower-variance policy gradients.
-The dense LLM-as-judge rewards replace sparse binary outcome signals with rich, fine-grained step-level supervision that directly evaluates reasoning quality, query quality, and answer correctness on a ternary scale.
-Our theoretical analysis provides formal guarantees on variance reduction, and our empirical results demonstrate significant improvements over both sparse-reward methods (Search-R1) and process-reward methods (StepSearch) across seven question answering benchmarks.
-Slate establishes a new state of the art for RL-based search-augmented reasoning, demonstrating that the combination of step-level exploration and dense model-based rewards is a powerful paradigm for training LLMs that reason effectively with external knowledge sources.
+We presented Slate, a training method for search-augmented reasoning whose key insight is that how step-level optimization is performed matters as much as what reward signal is used.
+By sampling $k$ continuations from a shared prefix, our method isolates variation to a single action with provable variance guarantees (up to $T$-fold reduction), a formal contribution absent from prior step-level reward approaches (Goldie et al., 2025; Wang et al., 2025).
+Combined with decomposed ternary rewards, Slate significantly outperforms both sparse-reward and process-reward baselines across seven benchmarks, with ablations confirming that truncated sampling yields gains above and beyond step-level rewards alone.
 Appendix A Appendix
@@
 […]
-20:  $\theta\leftarrow\theta+\alpha\nabla_{\theta}\Big[\frac{1}{k}\sum_{j=1}^{k}\mathcal{J}_{t}^{(j)}(\theta)-\beta\,\mathbb{D}_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\Big]$
+20:  Accumulate gradient: $\Delta\theta\mathrel{+}=\nabla_{\theta}\Big[\frac{1}{k}\sum_{j=1}^{k}\mathcal{J}_{t}^{(j)}(\theta)-\beta\,\mathbb{D}_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\Big]$
 21:   $\mathcal{J}_{t}^{(j)}(\theta)\!=\!\frac{1}{\sum_{l}I(y_{l})}\!\!\sum_{l:\,I(y_{l})=1}\!\!\min\!\big(\rho_{l}\hat{A}_{t}^{(j)},\;\text{clip}(\rho_{l},1{-}\epsilon,1{+}\epsilon)\hat{A}_{t}^{(j)}\big)$
@@
 30:end while
-31:end for
-Table 3: Main results (Exact Match) on Qwen2.5-3B-Base across seven QA benchmarks.
-Best results are bolded, second best are underlined.
-Search-R1 and Slate are trained on NQ+HotpotQA.
-StepSearch is trained on MuSiQue (19k) and does not report general QA results.
-Evaluated on Wiki-18 knowledge base.
-Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
-Direct 0.152 0.434 0.178 0.145 0.249 0.018 0.030 0.172
-CoT 0.168 0.451 0.196 0.163 0.257 0.032 0.072 0.191
-RAG 0.312 0.504 0.296 0.251 0.221 0.051 0.076 0.244
-IRCoT 0.286 0.478 0.268 0.171 0.196 0.059 0.245 0.243
-Search-o1 0.267 0.465 0.254 0.240 0.207 0.045 0.316 0.256
-SFT 0.301 0.491 0.286 0.191 0.248 0.039 0.112 0.238
-R1 (no search) 0.258 0.467 0.243 0.191 0.277 0.051 0.232 0.246
-ZeroSearch 0.348 0.531 0.365 0.260 0.234 0.056 0.096 0.270
-ReSearch 0.339 0.518 0.342 0.261 0.228 0.074 0.184 0.278
-Search-R1 0.406 0.587 0.435 0.284 0.273 0.049 0.088 0.303
-StepSearch 0.329 0.339 0.181 0.328
-Slate (Ours) 0.425 0.603 0.451 0.351 0.368 0.216 0.361 0.396
+[…]
+32:end for
+Note: Gradients are accumulated across all steps within a trajectory and applied in a single batched update after the trajectory completes.
+In practice, updates are further batched across multiple examples in the training batch.
 A.2 Gradient Estimators
@@
 Part 1 shows the truncated estimator is never worse; we now show it can be $T$ times better.
-The intuition is that the total trajectory variance $\text{Var}[R(\tau)]$ is the sum of variances from all $T$ steps, but the truncated estimator only “sees” the variance from one step—hence the $1/T$ factor.
-Under Assumption 2 (conditional independence), for each step $t$, the reward $r_{t}$ depends only on $a_{t}$ and $\tau_{<t}$, and future actions do not affect past rewards.
-Under this condition, the trajectory reward variance decomposes as:
-$\text{Var}[R(\tau)]$ $=\text{Var}\!\left[\sum_{t=1}^{T}r_{t}\right]=\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right]+\text{inter-step terms}.$
-Under Assumption 2, the inter-step covariance terms are non-negative: conditional independence ensures that $\text{Cov}(r_{t},r_{t^{\prime}}\mid\tau_{<t})\geq 0$ for $t<t^{\prime}$, since any remaining dependence flows only through the sequential prefix structure.
-Therefore:
+The intuition is that the total trajectory variance $\text{Var}[R(\tau)]$ is the sum of variances from all $T$ steps, but the truncated estimator only “sees” the variance from one step, hence the $1/T$ factor.
+Under Assumption 2 (conditional independence), for each step $t$, the reward $r_{t}$ depends only on $a_{t}$ and $\tau_{<t}$, and given the prefix $\tau_{<t}$, the step reward $r_{t}$ is independent of rewards at other steps conditioned on their respective prefixes.
+We establish the desired inequality by applying the law of total variance recursively.
+For any step $t$, the law of total variance gives:
+$\text{Var}[R(\tau)]=\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right]+\text{Var}_{\tau_{<t}}\!\left[\mathbb{E}[R(\tau)\mid\tau_{<t}]\right]\geq\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right].$ (18)
+From Part 1 (using Assumption 1), we already have $\text{Var}[r_{t}\mid\tau_{<t}]\leq\text{Var}[R(\tau)\mid\tau_{<t}]$ for each prefix $\tau_{<t}$. Summing over all steps:
+$\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right]\leq\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[R(\tau)\mid\tau_{<t}]\right].$ (19)
+Under Assumption 2, the per-step conditional variances account for distinct, non-overlapping sources of randomness (each step’s action is the sole source of variation given its prefix). Therefore:
 $\text{Var}[R(\tau)]\geq\sum_{t=1}^{T}\mathbb{E}_{\tau_{<t}}\!\left[\text{Var}[r_{t}\mid\tau_{<t}]\right].$ (20)
@@
 $=\frac{1}{T}\cdot\text{Var}[\hat{A}_{i}].$ (22)
-A.4 Sample Efficiency
+A.4 Proof Remarks
+Remark 1 (Bias-Variance Trade-off in the Estimator Comparison).
+Theorem 1 compares the variance of two different advantage estimators: the trajectory-level $\hat{A}_{i}$ (based on $R(\tau)$) and the step-level $\hat{A}_{t}^{(j)}$ (based on $r_{t}$).
+Because the truncated estimator targets the step-level reward rather than the full return, it is not an unbiased estimator of the trajectory-level advantage.
+The practical validity of this substitution rests on the step-level reward being a low-bias proxy for the contribution of action $a_{t}$ to the trajectory outcome, a condition favored by our setting’s short horizons ($T\leq 4$), externally grounded evaluation (retrieved documents are directly observable), and evaluative (rather than predictive) reward design (Section 3.3).
+We discuss this trade-off further in Appendix A.8.
+Remark 2 (Credit Assignment).
+The reward design provides an orthogonal benefit.
+By the data processing inequality, a binary EM reward provides at most 1 bit about the joint outcome of all $T$ actions, severely diluting the signal for any individual step.
+In contrast, the step-level LLM-judge reward $r_{t}$ directly evaluates $a_{t}$ in context, providing a substantially richer signal.
+Our ablation (Table 2) confirms that removing dense rewards causes larger drops than removing truncated sampling.
+We discuss the bias-variance trade-off of LLM-judge rewards in Appendix A.8.
+A.5 Sample Efficiency
 Proposition 2 (Sample Efficiency).
 To achieve the same advantage variance as standard GRPO with $G$ full-trajectory samples, the truncated step-level method requires only $G/T$ samples per step under the conditions of Theorem 1, yielding a $T$-fold reduction in total token generation cost.
-A.5 Proof of Proposition 2
+A.6 Proof of Proposition 2
 Proof.
@@
 In standard GRPO, we sample $G$ complete trajectories of average length $L$, costing $G\cdot L$ tokens.
-In our method, each truncated sample generates only one step’s worth of tokens—approximately $L/T$ tokens, since the full trajectory’s $L$ tokens are spread over $T$ steps—and we repeat this sampling at each of the $T$ steps.
+In our method, each truncated sample generates only one step’s worth of tokens, approximately $L/T$ tokens, since the full trajectory’s $L$ tokens are spread over $T$ steps, and we repeat this sampling at each of the $T$ steps.
 The total cost is therefore:
 […]
-[…]
-A.6 Variance Reduction and Convergence
+The two sources of savings, $T\times$ fewer samples needed (due to lower per-sample variance) and $T\times$ fewer tokens per sample (due to truncation), together yield a $T^{2}$ reduction; paying back one factor of $T$ for repeating across all $T$ steps gives a net $T$-fold improvement over the $G\cdot L$ tokens required by standard GRPO.
+∎
+A.7 Variance Reduction and Convergence
 Theorem 1 establishes that the truncated estimator yields lower-variance scalar advantages $\hat{A}_{t}^{(j)}$.
@@
 Beyond faster convergence, lower variance can also lead to better final solutions: cleaner gradients allow the optimizer to reliably descend into sharper, higher-performing regions of the loss landscape that noisy updates would overshoot or bounce out of, and they ensure that more of the finite training budget contributes useful learning signal rather than noise.
-A.7 Bias-Variance Trade-off
+A.8 Bias-Variance Trade-off
 Remark 3 (Bias-Variance Trade-off).
@@
 This aligns with the classical bias-variance trade-off in policy gradient methods, where moderate bias with substantially reduced variance yields better optimization dynamics (Schulman et al., 2015).
-A.8 Early-Termination Bonus Details
+A.9 Limitations
+Our truncated sampling constructs a single trajectory by greedily extending one selected action per step, which limits exploration compared to maintaining multiple full trajectories.
+While the short horizons ($T\leq 4$) and externally grounded rewards in our setting mitigate this, the approach may be less effective in domains with longer horizons or where locally promising actions frequently lead to dead ends.
+Additionally, our method relies on a large LLM judge (Gemma3-27B) for dense supervision, which introduces computational overhead and makes the overall improvement dependent on the quality of the judge model.
+A.10 Early-Termination Bonus Details
 The bonus term $\lambda\cdot(B-t)/B$ in Eq.
 6 encourages the model to produce an answer as soon as it has gathered sufficient information.
-Without such a term, the model may learn to issue superfluous search queries—each receiving a neutral or mildly positive reward from the LLM judge—even when the information needed to answer has already been retrieved.
+Without such a term, the model may learn to issue superfluous search queries, each receiving a neutral or mildly positive reward from the LLM judge, even when the information needed to answer has already been retrieved.
 The bonus is largest when the model answers early (e.g., $\lambda\cdot\frac{3}{4}$ at step t=1t{=}1 with B=4B{=}4) and zero when it exhausts the full budget (t=Bt{=}B), creating a progressive incentive to terminate sooner.
@@
 This ensures the policy gradient directly reinforces early termination when additional search steps are unlikely to improve the answer.
-A.9 Effect of Group Size $k$
+A.11 Why Process Rewards Succeed in Search but Not Math
+While process reward models have proven effective as verifiers for mathematical reasoning (Lightman et al., 2023), using step-level rewards as RL training signals has generally underperformed outcome-based verifiable rewards for policy optimization in math (Uesato et al., 2022).
+We argue that search-augmented reasoning possesses structural properties that make process rewards fundamentally more reliable as RL training signals, and that our truncated sampling design addresses the remaining risks.
+External Grounding Eliminates Capability Mismatch.
+In math, a process reward model (PRM) may assign high reward to an elegant proof strategy that the actor model cannot complete, for example, a sophisticated algebraic manipulation that only a larger model could carry out.
+This “capability mismatch” means the PRM rewards steps that look locally promising but lead to dead ends.
+In search, the “hard part” of each step is outsourced to the search engine: the model only needs to formulate a well-targeted query, and the retrieval system handles the actual information lookup.
+The capability ceiling for writing a good query is much lower than for completing an advanced proof, so the process reward cannot favor steps that exceed the actor’s abilities.
+Evaluative vs. Predictive Rewards.
+[…]
+Short Horizons and Natural Decomposability.
+Search trajectories in our setting have $T\leq 4$ steps, each consisting of a discrete think-then-query action with a natural boundary (the search engine call).
+Math proofs can involve dozens of tightly coupled steps where the value of step 5 depends critically on steps 15–20.
+With shorter, more modular trajectories, the gap between local step quality and trajectory-level outcome is inherently smaller, reducing the risk that process rewards mislead the policy.
+Truncated Sampling Prevents Local Minima.
+Even granting that search rewards are more locally informative, standard full-trajectory sampling still suffers from prefix confounding: a locally good step on a bad trajectory gets penalized, and a mediocre step on a lucky trajectory gets rewarded.
+This is exactly how local minima arise in PRM-based RL, the policy gets “trapped” by rewarding steps that merely co-occur with good outcomes.
+Our truncated sampling eliminates this confound: all $k$ candidates share the same prefix $\tau_{<t}$, so advantages reflect only the current action’s quality.
+The reward-weighted sampling strategy (temperature $\eta=0.7$) further prevents greedy collapse into local optima during trajectory construction.
+The ablation in Table 2 provides direct evidence: variant (a), which uses LLM-judge rewards without truncated sampling, underperforms full Slate most on the hardest benchmarks, precisely where prefix-level confounds are most harmful.
+Table 3: Main results (Exact Match) on Qwen2.5-3B-Base across seven QA benchmarks.
+Best results are bolded, second best are underlined.
+Search-R1 and Slate are trained on NQ+HotpotQA.
+StepSearch is trained on MuSiQue (19k) and does not report general QA results.
+Evaluated on Wiki-18 knowledge base.
+Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
+Direct 0.152 0.434 0.178 0.145 0.249 0.018 0.030 0.172
+CoT 0.168 0.451 0.196 0.163 0.257 0.032 0.072 0.191
+RAG 0.312 0.504 0.296 0.251 0.221 0.051 0.076 0.244
+IRCoT 0.286 0.478 0.268 0.171 0.196 0.059 0.245 0.243
+Search-o1 0.267 0.465 0.254 0.240 0.207 0.045 0.316 0.256
+SFT 0.301 0.491 0.286 0.191 0.248 0.039 0.112 0.238
+R1 (no search) 0.258 0.467 0.243 0.191 0.277 0.051 0.232 0.246
+ZeroSearch 0.348 0.531 0.365 0.260 0.234 0.056 0.096 0.270
+ReSearch 0.339 0.518 0.342 0.261 0.228 0.074 0.184 0.278
+Search-R1 0.406 0.587 0.435 0.284 0.273 0.049 0.088 0.303
+StepSearch 0.329 0.339 0.181 0.328
+Slate (Ours) 0.425 0.603 0.451 0.351 0.368 0.216 0.361 0.396
+A.12 Effect of Group Size $k$
 We study the impact of the number of truncated samples $k\in\{1,3,5,7\}$ per step on Qwen2.5-7B-Base (Table 4).
@@
 $k=7$ 0.449 0.411 0.244 0.491 0.399
-A.10 LLM-as-Judge Reward Prompts
+A.13 LLM-as-Judge Reward Prompts
 We provide the exact prompts used for the three LLM-as-judge reward components described in Section 3.3.
-In each prompt, the placeholders in curly braces (e.g., {context}, {thinking}) are filled with the corresponding trajectory content at evaluation time.
+In each prompt, the placeholders in curly braces are filled with the corresponding trajectory content at evaluation time.
 The judge is instructed to produce a chain-of-thought explanation before the score, enclosed in XML-style tags.